

********************************************************************
UARS WEEKLY STATUS REPORT - 04 JUNE 1999
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM gathered its normal solar, offset, and occultation data without
apparent problems.
We are evaluating the recently reprocessed V20 products before their
release to the DAAC.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
MLS was reactivated successfully on May 11 and operated until May 19 13:27 UTC
when the FOT needed to turn MLS off due to power constraints. MLS was reactived
on May 24 and operated through May 28.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
11 days CLA (7.0) PROD (L1-L3) (53R-36C-4F=13L)
48 days WIN (5.11) PROD (L1-L3) (1540R-632-4F=904L).
REHOST- Multinet is installed on the cluster allowing telnet and tcp/ip
capabilities. Optimizing Filetek primary files indices to improve performance
for Filetek data migration. Defined required logicals affecting Ingres
functions.
The CDHF will load April 1999 current day DAAC data in June 1999.
CDHF pending software upgrades: DEC AXP Fortran 6.5, DEC AXP OPEN
VMS 6.2, OpenIngres
***********
FOT
***********
REPORTING PERIOD
This report covers 22 May 1999 (Orbit 42039, GMT Day 142) through 28 May
1999 (Orbit 42143, GMT Day 148).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 143/1525 42063 Day Only
MLS 144/1707 42079 On
MLS 148/2344 42143 Off
The beta angle ranged between the angles of 77.1 degrees to 56.5 degrees
for this report period. The beta angle is now decreasing toward a minimum
angle of 0 degrees on 11 June 1999 (DOY 162).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 0.7 to 1.3 degrees C. Additional battery performance data
for each battery on May 21 1999 and May 28 1999 are provided in Appendix D.
The clock error ranged from 12.8 to -10.8 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = 00
% OUT-OF-LIMITS = 00
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 04
% MLS 04
% PEM 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 01
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 99/117 - 2287.496987
Transponder B: 99/118 - 2287.498739
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly Reports
are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 99/119 show the observatory data loss to be
24 hours, 13 minutes, 02.420 seconds (an increase of 12.096 seconds since
the last report period). This is a 0.0360 percent data loss which equals a
99.9640 percent data capture for the mission.
The increase of 12.096 seconds occurred over four (4) days:
7 0.416 seconds on Day 141 due to a data hit,
7 2.080 seconds on Day 142 due to no overlap,
7 0.448 seconds on Day 145 due to data hit,
7 9.152 seconds on Day 146 due to no overlap.
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
FOT personnel are supporting the IMOC (SCS-21) development effort. The
current work is a continuation of the SCS-21 UARS implementation effort
done last year which culminated in a level 1 demonstration.
The FOT continues to investigate the feasibility of a new PMON action to
change battery V/T curves based on loading to mitigate battery risks in sun
pointing safehold.
YEAR 2000 OPERABILITY SUPPORT
The end-to-end testing has been delayed until a new FORMATS software
release can be made to accommodate the special Y2K planning aids.
UARS PROPULSION MODULE B/D THRUSTER LINE FILL - No change
MEETINGS
FOT attended the weekly IMOC meetings on 26 May 1999.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver (forward-to-reverse) is scheduled for 11
June 1999 (DOY 162).
PERSONNEL CHANGES
None
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 28 May 1999
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
42060 143/1014 99-183 OBC Clk Adj 20.6 to 41.2 msec/day
42063 143/1525 99-184 PEM Power ON, Day Only Ops
42079 144/1707 99-162 MLS Power ON
42100 146/0249 99-185 OBC Clk Adj 41.2 to 20.6 msec/day
42109 146/1646 99-186 MLS Scan Rate Adj.
42121 147/1131 99-187 OBC Clk Adj 20.6 to 0.00 msec/day
42139 148/1727 99-188 MLS MLS RESET
42142 148/2202 99-189 OBC Clk Adj 0.00 to 20.6 msec/day
42143 148/2344 99-190 MLS Power OFF
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 28 May 1999
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
_______________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 28 May 1999
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
42060 143/10:18 (23 May 99) from 20.6 to 41.2 msec/day
42100 146/02:56 (25 May 99) from 41.2 to 20.6 msec/day
42121 147/11:38 (27 May 99) from 20.6 to 00.0 msec/day
42142 148/22:05 (28 May 99) from 00.0 to 20.6 msec/day
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 28 May 1999
21 May 1999 - GMT Day 141 Beta = 77.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -89.6/-106.4 -2.11/-3.05 +0.0/ +0.0 16.5 0.000 0.0
2 +11.2/-11.2 +2.92/+1.34 +15.2/-16.0 25.1 1.077 12.8
3 +22.4/-44.8 +2.29/+0.71 +15.2/-16.0 25.1 1.081 12.6
28 May 1999 - GMT Day 148 Beta = 56.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 -89.6/-112.0 -3.68/-4.94 +0.0/ +0.0 16.5 0.000 0.0
2 +22.4/-11.2 +5.15/+3.87 +25.6/-16.0 24.8 1.022 21.0
3 +33.6/-22.4 +4.51/+2.60 +24.8/-16.0 24.8 1.026 20.6
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes V/T 5
APPENDIX E
UARS ANOMALY REPORTS
Week ending 28 May 1999
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-027 FOT 42100 146/0249 20905 I channel dropout/no
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
None this report period
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
99-023 FOT see list May 99 14170 Generic Late Acq
99-027 FOT 42100 146/0249 20905 I channel dropout/no
SEVNTFW99/22